# Phenomenological study of the angle between jet axes in heavy-ion collisions

### v1

@@ Abstract
This paper presents a phenomenological study on the angle between the Standard and the Winner-Take-All (WTA) jet axes ($\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$) in high-energy nuclear collisions. The $p$+$p$ baseline is provided by the Pythia8 event generator. The in-medium jet propagation is simulated by the linear Boltzmann transport (Lbt) model, which considers both the elastic and inelastic jet-medium interactions. Our theoretical results calculated by the Lbt model show that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ is narrower than that in $p$+$p$, which agrees well with the recent ALICE measurements. The narrowing of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ seems to violate the $p_{T}$-broadening nature of the jet quenching effect, usually explained by the influence of “selection bias”. However, the physical details still need to be fully understood. Utilizing a matching-jet method to track the jet evolution in the QGP to remove the selection bias in the Monte Carlo simulations, we observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution becomes broader due to the jet-medium interactions. At the same time, by rescaling the quark/gluon-jet fractions in Pb+Pb collisions to be the same as that in $p$+$p$, we find that the fraction change may not significantly influence the modification pattern of jet $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$. On the other hand, the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline, and such a biased comparison between $p$+$p$ and A+A conceals the actual jet-broadening effect in the experimental measurements. The investigations presented in this paper will deepen our understanding of the relationship between the actual intra-jet modifications in the QGP and the experimental observations.
@@ I Introduction
Exploring the properties of the quark-gluon plasma (QGP), a new state of nuclear matter formed at extremely hot and dense conditions, is one of the most important motivations of the heavy-ion collision program at the RHIC and the LHC Freedman and McLerran (1977); Shuryak (1978); Connors et al. (2018); Blaizot and Mehtar-Tani (2015); Braun-Munzinger et al. (2016); Cunqueiro and Sickles (2022); Cao and Qin (2023); Apolinário et al. (2022). The high $p_{T}$ parton or jets generated at a very early stage of the collisions will strongly interact with the medium and dissipate energy when traversing through the QGP, called the “jet quenching” phenomenon. Therefore, jets are widely regarded to be an excellent probe for unveiling the mystery of such deconfined quark soup Wang and Gyulassy (1992); Gyulassy et al. (2004); Mehtar-Tani et al. (2013); Qin and Wang (2015); Wang (1998); Wang and Wang (2001, 2002); Vitev and Zhang (2010); Neufeld et al. (2011); He et al. (2020); Acharya et al. (2023); Kang et al. (2023); Dai et al. (2013); Yang et al. (2023a); Wang et al. (2023a); Xie et al. (2023); Yang et al. (2023b). A series of recently developed jet substructure observables have been used to perform the investigations for further understanding the QGP medium and nucleon structure, including the scaled groomed jet radius $\theta_{g}$, the groomed jet momentum splitting fraction $z_{g}$ Ringer et al. (2020); Acharya et al. (2020, 2022); Wang et al. (2023b); Aad et al. (2023); Zhang et al. (2023); Tachibana et al. (2023), the number of “Soft Drop” branches $n_{\mathrm{SD}}$ Acharya et al. (2020), the radial profile Li et al. (2023); Wang et al. (2021a, b, 2019), the generalised angularities $\lambda_{\beta}^{\kappa}$ Ehlers (2022), the jet charge $Q^{\kappa}$ Chen et al. (2020); Sirunyan et al. (2020), the groomed jet mass Sirunyan et al. (2018), the splitting parameter $\sqrt{d_{12}}$, the angular separation $\Delta R_{12}$ ATLAS Collaboration (2023), the $N$-subjettiness, and the sub-jet fragmentation Mulligan (2022), etc.
@@
Besides the observables mentioned above, there is a fascinating observable—the angle between differently defined jet axes, $\Delta R_{\mathrm{axis}}$, which probes a vast phase space of the jet formation and evolution ALICE Collaboration (2022). The concept of the angle between jet axes $\Delta R_{\mathrm{axis}}$ was firstly proposed in Ref. Cal et al. (2020), in which $\Delta R_{\mathrm{axis}}$ was calculated at next-to-leading logarithmic accuracy and consistent with the simulations of the Monte Carlo event generators. Subsequently, the ALICE collaboration reported experiment measurements of the angle between jet axes in $p$+$p$ and Pb+Pb collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ ALICE Collaboration (2022); Acharya et al. (2023).
@@
The existing studies show that the physical meanings of the $\Delta R_{\mathrm{axis}}$ observables are clear and straightforward. $\Delta R_{\mathrm{axis}}$ measures the geometric distance of two different axes determined by two different algorithms for the same reconstructed jet in the $(\eta,\phi)$ plane. Various algorithms determine different jet axes, each with varying degrees of sensitivity to the soft radiation. More specifically, the “Soft Drop” grooming algorithm removes the soft wide-angle radiation in the jets, and the degree of this grooming can be adjusted by parameters Larkoski et al. (2014a). The WTA axis is obtained by reclustering using the “Winner-Take-All” recombination scheme, which tracks the energetic collinear radiation. The WTA axis is typically aligned with the most energetic components of the jet, and the effect of soft radiation is power suppressed in the WTA axis ALICE Collaboration (2022); Cal et al. (2020). Systematic discussions of $\Delta R_{\mathrm{axis}}$ in $p$+$p$ collisions are performed in Ref. Cal et al. (2020).
@@
In nucleus-nucleus collisions, due to the different sensitivity to the soft radiation for different definitions of the jet axis, the $\Delta R_{\mathrm{axis}}$ provides a unique opportunity to gain insight into the energy loss mechanisms of jet-medium interactions. Recently the ALCIE collaboration has reported the first measurement on the $\Delta R_{\mathrm{axis}}$ distributions of charged-particle jets in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ Acharya et al. (2023). It is found that $\Delta R_{\mathrm{axis}}$ has narrower distributions in Pb+Pb collisions than in $p$+$p$. The experimental results show different modification patterns compared to the theoretical calculations with an intra-jet $p_{T}$-broadening Ringer et al. (2020). In Ref. Acharya et al. (2023), it is implied that the “selection bias” may play an essential role during the jet finding in A+A collisions. In other words, the minimum $p_{T}^{\rm jet}$ threshold biases the jet selection to favor the candidates with less energy loss Renk (2013); Cunqueiro and Sickles (2022), but the details in physics are unclear. Moreover, due to the different Casimir color factors ($C_{F}=4/3$ for quark and $C_{A}=3$ for gluon), the gluon-initiated jets will lose more energy than quark-initiated jets during medium-induced interactions, which leads to a reduced gluon-jet fraction of jet sample in A+A collision Chen et al. (2020); Li and Vitev (2020); Sirunyan et al. (2020). Since the gluon-jet is expected to interact more with the QGP medium, the reduced gluon-jet fractions and relatively enhanced quark-jet fraction are believed to be the main reason that narrows the $\Delta R_{\mathrm{axis}}$ distribution in Pb+Pb collisions compared to the vacuum case Du et al. (2020); Brewer et al. (2022). Further detailed theoretical explorations are necessary to disentangle the influences of the selection bias and the quark/gluon-jet fraction changes to the observation in the experiment, which is essential to understand the relationship between the actual intra-jet modifications in the QGP and the experimental observations.
@@
This paper presents a theoretical study on the angle between the Standard and the WTA jet axes ($\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$) in high-energy nuclear collisions. The $p$+$p$ baseline is provided by the Pythia8 event generator, and the in-medium jet propagation is simulated by the linear Boltzmann transport (Lbt) model. We will perform the calculations of the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ compared to the recently reported ALICE measurements. Furthermore, we will investigate the influences of the event “selection bias” and the quark/gluon-jet fraction changes to the modification patterns of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in nucleus-nucleus collisions. We will use the “matching-jet” method to track the jet evolution in the QGP, in which the realistic jet modification can be unveiled. In addition, we will perform a detailed analysis to demonstrate that the critical factor concealing such jet-broadening effects in experimental measurement is the biased comparison between two jet samples in $p$+$p$ and Pb+Pb collisions. At the same time, the influence of the quark/gluon-jet fraction changes is limited.
@@ II Framework
When employing different reclustering/grooming algorithms to process a given jet, we will get some new jets slightly different from the original one. $\Delta R_{\mathrm{axis}}$ is a novel jet-substructure observable proposed in Ref. Cal et al. (2020) to study the angular distance between the two jet axes defined with different jet algorithms:
$\Delta R_{\rm axis}^{a-b}=\sqrt{(y_{\rm axis}^{a}-y_{\rm axis}^{b})^{2}+(\phi_{\rm axis}^{a}-\phi_{\rm axis}^{b})^{2}}$ (1)
where $a$ and $b$ denote any two different axis algorithms, $y$ and $\phi$ are the rapidity and the azimuthal angle. According to different reclustering/grooming algorithms, three types of $\Delta R_{\mathrm{axis}}$ combination are extensively investigated in jet physics at the LHC Cal et al. (2020); ALICE Collaboration (2022); Ehlers (2022); Acharya et al. (2023): WTA – Standard, WTA – SoftDrop and Standard – SoftDrop. This work will focus on the angle between the standard and WTA axes.
At the LHC, almost universally employed the $\mathrm{anti\mbox{-}}k_{t}$ algorithm Cacciari et al. (2008) to reconstruct jets Gauld et al. (2023), so we call the jet axis obtained using the $\mathrm{anti\mbox{-}}k_{t}$ algorithm (with default $E$ recombination scheme) clustering as the “Standard” axis. The WTA is a recombination scheme that specifies how to combine the momenta when merging two subbranches during the clustering procedure. Specifically, we directly assign the transverse momentum sum of the two subbranches as the transverse momentum of the merged branch and designate the direction of the harder of two subbranches as the direction of the merged branch ALICE Collaboration (2022); Bertolini et al. (2014),
@@
where the subscript $r$ denotes the merged branch; $i$ and $j$ denote the two prongs, respectively. The WTA scheme is infrared/collinear safe and can powerfully suppress the effect of soft radiation Cal et al. (2020); ALICE Collaboration (2022). Because of these features, the WTA scheme has been used to study jet-like event shapes Bertolini et al. (2014), recoil-free observables Larkoski et al. (2014b); Neill et al. (2019), flavor identification of jets Caletti et al. (2022), etc. Back to our work, we recluster the constituents of the “Standard” jet using the C/A algorithm Dokshitzer et al. (1997); Wobisch and Wengler (1998) with the WTA $p_{T}$ recombination scheme, and then we can get the WTA axis.
A Monte Carlo event generator Pythia v8.309 Bierlich et al. (2022) with Monash 2013 Skands et al. (2014) tune is used to simulate jet productions in $p$+$p$ collisions as the baseline. We use charged-particles with $p_{T}>0.15~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and pseudorapidity range $|\eta|<0.9$ to reconstruct jets. All jets are reconstructed by the $\mathrm{anti\mbox{-}}k_{t}$ algorithm with the $E$ recombination scheme and the distance parameter $R=0.2$ using FastJet v3.4.0 package Cacciari and Salam (2006); Cacciari et al. (2012).
We employ the linear Boltzmann transport (Lbt) model Wang and Zhu (2013); He et al. (2015); Cao et al. (2016, 2018) to simulate both elastic and inelastic scattering processes of the jet shower partons and the thermal recoil partons in the QGP medium. The Lbt model performed well in a series of jet quenching measurements, which was highly consistent with the experimental results Wang and Zhu (2013); Cao et al. (2016, 2018); Luo et al. (2018); Zhang et al. (2018). In the Lbt model, the linear Boltzmann transport equation is used to describe the elastic scattering process,
@@
where $f_{i}(p_{i})(i=1,2,3,4)$ are the phase-space distribution of partons that participate in the reaction; $\left|\mathcal{M}_{12\to 34}\right|^{2}$ is the leading-order (LO) elastic scattering matrix elements Combridge (1979); Eichten et al. (1984); in order to avoid possible collinear divergence of $\left|\mathcal{M}_{12\to 34}\right|^{2}$ for massless partons, here, a Lorentz-invariant regularization condition $S_{2}(\hat{s},\hat{t},\hat{u})\equiv\theta\left(\hat{s}\geqslant 2\mu_{D}^{2}\right)\theta\left(-\hat{s}+\mu_{D}^{2}\leqslant\hat{t}\leqslant-\mu_{D}^{2}\right)$ is introduced Auvinen et al. (2010); $\mu_{D}^{2}$ is the Debye screening mass. It should be noted that the medium response effect is included in the Lbt model during our simulations. The higher-twist approach Guo and Wang (2000); Wang and Guo (2001); Zhang and Wang (2003); Zhang et al. (2004, 2005) describes the inelastic scattering of medium-induced gluon radiation,
@@
where $x$ is the energy fraction of the emitted gluon taken from its parent parton with mass $m$, $k_{\perp}$ is the transverse momentum of the radiated gluon, $P(x)$ is the splitting function, $\hat{q}$ is the jet transport coefficient (specific details can be found in Ref He et al. (2015)), and $\tau_{f}$ is the formation time of the radiated gluons in the QGP medium. The background hydrodynamic profile for the bulk QGP medium is provided by simulations from the 3$+$1D Music relativistic second-order viscous hydro model Schenke et al. (2010, 2011); Paquet et al. (2016) with average initial conditions from the 3D TrENTo model Moreland et al. (2015); Ke et al. (2017).
After the evolutions of the parton shower in the medium are completed, we need to perform hadronization to obtain the hadron-level events that can be observed in the experiment. In this work, we first construct strings using the colorless method from the JetScape framework based on minimization criteria Putschke et al. (2019) and then perform hadronization and hadron decays using the Lund string model Bierlich et al. (2022); Andersson et al. (1983); Sjostrand (1984) provided by Pythia. The JetScape collaboration has already demonstrated in works Kumar et al. (2023); Tachibana et al. (2023) that this model is reliable.

### v2

@@ Abstract
This paper presents a phenomenological study on the angle between the Standard and Winner-Take-All (WTA) jet axes ($\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$) in high-energy nuclear collisions. The $p$+$p$ baseline is provided by the Pythia8 event generator. The in-medium jet propagation is simulated by the linear Boltzmann transport (Lbt) model, which considers both the elastic and inelastic jet-medium interactions. Our theoretical results calculated by the Lbt model show that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ is narrower than that in $p$+$p$, which agrees well with the recent ALICE measurements. The narrowing of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$, which seems to violate the nature of intra-jet broadening due to jet quenching, may be attributed to the influence of “selection bias”. However, the physical details still need to be fully understood. Utilizing a matching-jet method to track the jet evolution in the QGP to remove the selection bias in the Monte Carlo simulations, we observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution becomes broader due to the jet-medium interactions. At the same time, by rescaling the quark/gluon-jet fractions in Pb+Pb collisions to be the same as that in $p$+$p$, we find that the fraction change may not significantly influence the modification pattern of jet $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$. On the other hand, the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline, and such a biased comparison between $p$+$p$ and A+A conceals the actual intra-jet broadening effect in the experimental measurements. The investigations presented in this paper will deepen our understanding of the relationship between the actual intra-jet modifications in the QGP and the experimental observations.
@@ I Introduction
Exploring the properties of the quark-gluon plasma (QGP), a new state of nuclear matter formed at extremely hot and dense conditions, is one of the most important motivations of the heavy-ion collision program at the RHIC and the LHC [1, 2, 3, 4, 5, 6, 7, 8]. The high $p_{\mathrm{T}}$ parton or jets generated at a very early stage of the collisions will strongly interact with the medium and dissipate energy when traversing through the QGP, called the “jet quenching” phenomenon. Therefore, jets are widely regarded to be an excellent probe for unveiling the mystery of such deconfined quark soup [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. A series of recently developed jet substructure observables have been used to perform the investigations for further understanding the QGP medium and nucleon structure, including the scaled groomed jet radius $\theta_{g}$, the groomed jet momentum splitting fraction $z_{g}$ [26, 27, 28, 29, 30, 31, 32], the number of “Soft Drop” branches $n_{\mathrm{SD}}$ [27], the radial profile [33, 34, 35, 36], the generalised angularities $\lambda_{\beta}^{\kappa}$ [37], the jet charge $Q^{\kappa}$ [38, 39], the groomed jet mass [40], the splitting parameter $\sqrt{d_{12}}$, the angular separation $\Delta R_{12}$ [41], the $N$-subjettiness, and the sub-jet fragmentation [42], etc.
@@
Besides the observables mentioned above, there is a fascinating observable—the angle between differently defined jet axes, $\Delta R_{\mathrm{axis}}$, which probes a vast phase space of the jet formation and evolution [43]. The concept of the angle between jet axes $\Delta R_{\mathrm{axis}}$ was firstly proposed in Ref. [44], in which $\Delta R_{\mathrm{axis}}$ was calculated at next-to-leading logarithmic accuracy and consistent with the simulations of the Monte Carlo event generators. Subsequently, the ALICE collaboration reported experiment measurements of the angle between jet axes in $p$+$p$ and Pb+Pb collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ [43, 19].
@@
The existing studies show that the physical meanings of the $\Delta R_{\mathrm{axis}}$ observables are clear and straightforward. $\Delta R_{\mathrm{axis}}$ measures the geometric distance of two different axes determined by two different algorithms for the same reconstructed jet in the $(\eta,\phi)$ plane. Various algorithms determine different jet axes, each with varying degrees of sensitivity to the soft radiation. More specifically, the “Soft Drop” grooming algorithm removes the soft wide-angle radiation in the jets, and the degree of this grooming can be adjusted by parameters [45]. The WTA axis is obtained by reclustering using the “Winner-Take-All” recombination scheme, which tracks the energetic collinear radiation. The WTA axis is typically aligned with the most energetic components of the jet, and the effect of soft radiation is power suppressed in the WTA axis [43, 44]. Systematic discussions of $\Delta R_{\mathrm{axis}}$ in $p$+$p$ collisions are performed in Ref. [44].
@@
In nucleus-nucleus collisions, due to the different sensitivity to the soft radiation for different definitions of the jet axis, the $\Delta R_{\mathrm{axis}}$ provides a unique opportunity to gain insight into the energy loss mechanisms of jet-medium interactions. Recently the ALICE collaboration has reported the first measurement on the $\Delta R_{\mathrm{axis}}$ distributions of charged-particle jets in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ [19]. It is found that $\Delta R_{\mathrm{axis}}$ has narrower distributions in Pb+Pb collisions than in $p$+$p$. The experimental results show different modification patterns compared to the theoretical calculations with an intra-jet broadening [26]. In Ref. [19], it is implied that the “selection bias” may play an essential role during the jet finding in A+A collisions. In other words, the minimum $p_{\mathrm{T}}^{\mathrm{jet}}$ threshold biases the jet selection to favor the candidates with less energy loss [46, 6], but the details in physics are unclear. Moreover, due to the different Casimir color factors ($C_{F}=4/3$ for quark and $C_{A}=3$ for gluon), the gluon-initiated jets will lose more energy than quark-initiated jets during medium-induced interactions, which leads to a reduced gluon-jet fraction of jet sample in A+A collisions [38, 47, 39]. Since the gluon-jet is expected to interact more with the QGP medium, Ref. [19] suggests that the reduced gluon-jet fraction (from an alternative perspective, a relatively enhanced quark-jet fraction) is believed to be the main reason for the narrower $\Delta R_{\mathrm{axis}}$ distribution in Pb+Pb collisions compared to the vacuum case. Further detailed theoretical explorations are necessary to disentangle the influences of the selection bias and the quark/gluon-jet fraction changes to the observation in the experiment, which is essential to understand the relationship between the actual intra-jet modifications in the QGP and the experimental observations.
@@
This paper presents a theoretical study on the angle between the Standard and WTA jet axes ($\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$) in high-energy nuclear collisions. The $p$+$p$ baseline is provided by the Pythia8 event generator, and the in-medium jet propagation is simulated by the linear Boltzmann transport (Lbt) model. We will perform the calculations of the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ compared to the recently reported ALICE measurements. Furthermore, we will investigate the influences of the event “selection bias” and the quark/gluon-jet fraction changes to the modification patterns of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in nucleus-nucleus collisions. We will use the “matching-jet” method to track the jet evolution in the QGP, in which the realistic jet modification can be unveiled. In addition, we will perform a detailed analysis to demonstrate that the critical factor concealing such intra-jet broadening effects in experimental measurement is the biased comparison between two jet samples in $p$+$p$ and Pb+Pb collisions. At the same time, the influence of the quark/gluon-jet fraction changes is limited.
@@ II Framework
When employing different reclustering/grooming algorithms to process a given jet, we will get some new jets slightly different from the original one. The $\Delta R_{\mathrm{axis}}$ is a novel infrared and collinear (IRC) safe jet-substructure observable proposed in Ref. [44] to study the angular distance between the two jet axes defined with different jet algorithms:
$\Delta R_{\rm axis}^{a-b}=\sqrt{(y_{\rm axis}^{a}-y_{\rm axis}^{b})^{2}+(\phi_{\rm axis}^{a}-\phi_{\rm axis}^{b})^{2}},$ (1)
where $a$ and $b$ denote any two different axis algorithms, $y$ and $\phi$ are the rapidity and the azimuthal angle. According to different reclustering/grooming algorithms, three types of $\Delta R_{\mathrm{axis}}$ combination are extensively investigated in jet physics at the LHC [44, 43, 37, 19]: WTA – Standard, WTA – SoftDrop and Standard – SoftDrop. This work will focus on the angle between the Standard and WTA axes.
At the LHC, almost universally employed the $\mathrm{anti\mbox{-}}k_{t}$ algorithm [48] to reconstruct jets [49], so we call the jet axis obtained using the $\mathrm{anti\mbox{-}}k_{t}$ algorithm (with default $E$ recombination scheme) clustering as the “Standard” axis. The WTA is a recombination scheme that specifies how to combine the momenta when merging two subbranches during the clustering procedure. Specifically, we directly assign the transverse momentum sum of the two subbranches as the transverse momentum of the merged branch and designate the direction of the harder of two subbranches as the direction of the merged branch [43, 50],
@@
where the subscript $r$ denotes the merged branch; $i$ and $j$ denote the two prongs, respectively. The WTA scheme is infrared/collinear safe and can powerfully suppress the effect of soft radiation [44, 43]. Because of these features, the WTA scheme has been used to study jet-like event shapes [50], recoil-free observables [51, 52], flavor identification of jets [53], etc. Back to our work, we recluster the constituents of the “Standard” jet using the C/A algorithm [54, 55] with the WTA $p_{\mathrm{T}}$ recombination scheme, and then we can get the WTA axis.
A Monte Carlo event generator Pythia v8.309 [56] with Monash 2013 [57] tune is used to simulate jet productions in $p$+$p$ collisions as the baseline. We use charged-particles with $p_{\mathrm{T}}>0.15~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and pseudorapidity range $|\eta|<0.9$ to reconstruct jets. All jets are reconstructed by the $\mathrm{anti\mbox{-}}k_{t}$ algorithm with the $E$ recombination scheme and the distance parameter $R=0.2$ using FastJet v3.4.0 package [58, 59].
We employ the linear Boltzmann transport (Lbt) model [60, 61, 62, 63] to simulate both elastic and inelastic scattering processes of the jet shower partons and the thermal recoil partons in the QGP medium. The Lbt model performed well in a series of jet quenching measurements, which was highly consistent with the experimental results [60, 62, 63, 64, 65]. In the Lbt model, the linear Boltzmann transport equation is used to describe the elastic scattering process,
@@
where $f_{i}(p_{i})(i=1,2,3,4)$ are the phase-space distribution of partons that participate in the reaction; $\left|\mathcal{M}_{12\to 34}\right|^{2}$ is the leading-order (LO) elastic scattering matrix elements [66, 67]; in order to avoid possible collinear divergence of $\left|\mathcal{M}_{12\to 34}\right|^{2}$ for massless partons, here, a Lorentz-invariant regularization condition $S_{2}(\hat{s},\hat{t},\hat{u})\equiv\theta\left(\hat{s}\geqslant 2\mu_{D}^{2}\right)\theta\left(-\hat{s}+\mu_{D}^{2}\leqslant\hat{t}\leqslant-\mu_{D}^{2}\right)$ is introduced [68]; $\mu_{D}^{2}$ is the Debye screening mass. It should be noted that the medium response effect is included in the Lbt model during our simulations. The higher-twist approach [69, 70, 71, 72, 73] describes the inelastic scattering of medium-induced gluon radiation,
@@
where $x$ is the energy fraction of the emitted gluon taken from its parent parton with mass $m$, $k_{\perp}$ is the transverse momentum of the radiated gluon, $P(x)$ is the splitting function, $\hat{q}$ is the jet transport coefficient (specific details can be found in Ref [61]), and $\tau_{f}$ is the formation time of the radiated gluons in the QGP medium. The background hydrodynamic profile for the bulk QGP medium is provided by simulations from the 3$+$1D Music relativistic second-order viscous hydro model [74, 75, 76] with the averaged initial condition from the 3D TrENTo model [77, 78]. The evolution of the background medium was carried out in a few key steps. First, we employed the 3D TrENTo model to generate about 2000 event-by-event initial conditions for the most central 0-10% of Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV. Next, we averaged these initial conditions to obtain the average initial state profile. Ignoring the evolution of the pre-equilibrium stage, we directly used the averaged initial-state condition as the input for the 3+1D Music model to simulate the evolution of the QGP medium. Consequently, our current framework does not include event-by-event fluctuations [79] or pre-equilibrium effects.
After the evolutions of the parton shower in the medium are completed, we need to perform hadronization to obtain the hadron-level events that can be observed in the experiment. In this work, we first construct strings using the colorless method from the JetScape framework based on minimization criteria [80] and then perform hadronization and hadron decays using the Lund string model [56, 81, 82] provided by Pythia. The JetScape collaboration has already demonstrated in works [83, 32] that this model is reliable.